The ABC Company has developed a new cellphone model. The engineering department claims that its battery lasts for 4 days. In order to test this claim, the company selects a random sample of 100 new cellphones so that this sample has a mean battery life of 2.5 days with a standard deviation of 1 day. 



1
Expert's answer
2022-06-06T23:08:50-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=4"

"H_1:\mu\not=4"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=99" and the critical value for a two-tailed test is "t_c =2.626405."

The rejection region for this two-tailed test is "R = \{t:|t|>2.626405\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{2.5-4}{1/\sqrt{100}}=-15"


Since it is observed that "|t|=15>2.626405=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=99" degrees of freedom, "t=-4" is "p=0," and since "p= 0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 4, at the "\alpha = 0.01" significance level.